OpenStudy (anonymous):

Solve: -3(9x + 8) = -x – 50

OpenStudy (anonymous):

-27x-24=-x-50 ===> 26x=26 ==> x=26/26 = 1

OpenStudy (anonymous):

x=1

OpenStudy (anonymous):

x=1